Assume that y varies directly as x. Find y when x = 54 if y = 1/4 when x = 3/2.
Answer by midwood_trail(310)

Assume that y varies directly as x. Find y when x = 54 if y = 1/4 when x = 3/2.
y varies directly as x is written y = x
This is a proportion.
(y/54) = (1/4)/(3/2)
Can you solve for y now?
If not, write back.